OpenStudy (anonymous):

Which is a rational function? a. y=√x-3 b. y=x²-3x+5 c. y=x-5/x d.y=5

OpenStudy (akshay_budhkar):

certainly y=5 looks rational

OpenStudy (anonymous):

from the word rational, "ratio", which among the options has a fraction? it's C

Directrix (directrix):

A rational function is defined as the quotient of two polynomial functions. Grouping symbols are important in mathematics because they provide clarity. While, c. y=x-5/x appears to be the answer, it is not clear if c) mean y = (x - 5)/x or y = x - (5/x) which could be converted to y = (x^2-5) / x. I suppose that either interpretation leads to a rational function although grouping symbols would dissolve the ambiguity here.

OpenStudy (anonymous):

@directrix. correct